OpenStudy (anonymous):

I'm studying for a test and I really need the help please. A rectangles width is 6 feet less than its length. If the area of the rectangle is 247 square feet, what is its length, in feet?

OpenStudy (anonymous):

I have some examples from my notes, but they didnt help me out as much as I had hoped

OpenStudy (anonymous):

so w=l-6

OpenStudy (anonymous):

the area is l x w

OpenStudy (anonymous):

so l(l-6) =274 which is equal to l^2-6l-274

OpenStudy (anonymous):

so use the quadratic formula and that will give you the answer

OpenStudy (anonymous):

okay, ill try it out.

OpenStudy (anonymous):

I got x=566

OpenStudy (anonymous):

\[(L-6) L=247 \]\[\{L\to -13,L\to 19\} \]\[L=19 \]

OpenStudy (anonymous):

247 is square feet not the length. how did you get that??

OpenStudy (anonymous):

width times length is area in square units.

OpenStudy (anonymous):

okay then.

OpenStudy (anonymous):

instead of the quad formula you could also factor l^2-6l-247

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

(L - 6) L == 247 // Solve // Flatten Is the actual text that was fed to Mathematica to solve for L .